What is the difference between parallel and coincide?
example: 18=18

Answer by longjonsilver(2297)

in terms of straight lines you mean?

y = x
y = x + 2
y = x + 34
y = x - 17

all these are parallel...same gradient (+1) but cross the y-axis and different places.

Now, y=x and 3y = 3x are the same equation...they are coincident. As are 2y = 3x - 6 and y = (3/2)x - 3.

Hope that helps?

jon.